OpenStudy (anonymous):

The area of the trapezium is 105 cm2 and its height is 7 cm. If one of the parallel side is longer than the other by 6 cm. Find the two parallel sides

OpenStudy (bradely):

A =(h/2)(a+b) 105=(7/2)(x+x+6) 30=2x+6 2x=24 x=12 a=12 b=18 Source: http://www.mathskey.com/question2answer/ask

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I have one more question A regular human hand contains 27 different bones. There are 206 bones altogether in the whole human body. Approximately what percent of the bones in a human are in one hand.

OpenStudy (bradely):

(27*100)/206 = 13% Source: http://www.mathskey.com/question2answer/ask

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

.In a Mathematics class , the teacher asked Ajay and Karan to find the value of 2x-y , If 5raise to x−4 =625 and 2 raise to 2y−2 =2 raise to 8. Ajay got the answer as 10 whereas Karan got 11as the answer. Whom do you think is right? Justify your answer by showing the suitable steps.

OpenStudy (anonymous):

I am confuse in this question plzz answer me

OpenStudy (bradely):

Please check the answer http://www.mathskey.com/question2answer/16795/maths

OpenStudy (anonymous):

thank you so much !!!!!

OpenStudy (anonymous):

The length of the diagonals of a rhombus are 24 cm and 18 cm.Find the perimeter and area of the rhombus.

OpenStudy (anonymous):

plzz help

OpenStudy (bradely):

Area =( 1/2 )d1d2 Substitute and find the area Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

actually i got area but i couldn't find perimeter

OpenStudy (anonymous):

i got the answer thank you

OpenStudy (anonymous):

[If 2^{x} \times 5^{x} =1000 Find x\]

OpenStudy (bradely):

please check the answer http://www.mathskey.com/question2answer/17180/maths-help